OpenStudy (anonymous):

Silly question and really easy one probably... what is the square root of 194??

OpenStudy (geometry_hater):

its 13.92

OpenStudy (anonymous):

THANK YOU!!!!!

OpenStudy (anonymous):

do you want an exact answer? if so 19.92 is not right?

OpenStudy (anonymous):

13.92 I mean

OpenStudy (anonymous):

Yes ... I have a question that is getting the best of me and the math is just not adding up to what the answer should be

OpenStudy (anonymous):

give the whole question?

OpenStudy (anonymous):

because \[\sqrt{194}\] is irrational number

OpenStudy (geometry_hater):

Square root of 194 is 13.93 actually i didn't round up

OpenStudy (anonymous):

13.92 is a approx answer

OpenStudy (anonymous):

Find the perimeter and the area of a right triangle if the shortest side is 5cm and the longest side is 13cm

OpenStudy (anonymous):

13.92838827718412 and on and on

OpenStudy (anonymous):

okay

OpenStudy (anonymous):



OpenStudy (anonymous):

okay use the Pythag Thereoem \[a ^{2} + b ^{2} = c ^{2}\]

OpenStudy (anonymous):

so oh wait it say 13 is the longest side

OpenStudy (anonymous):

right.... does it matter which side is AB or C or am i getting to detailed for no reason

OpenStudy (anonymous):

because the hypotenuses is always the longest side in a right triangle

OpenStudy (anonymous):

5`2 + b`2 = 13`2

OpenStudy (anonymous):

Ok, I am following you..

OpenStudy (anonymous):

25 + b`2 = 169 -25 -25 b`2 = 144 b = 12

OpenStudy (anonymous):

then add all the sides to get the perimeter

OpenStudy (anonymous):

sorry I have to go, I hope I helped

OpenStudy (anonymous):

You did thank you so much!

OpenStudy (anonymous):

:)